OpenStudy (anonymous):

Check my work please? What is the simplified form of the quantity 3 over x minus 2 over y all over the quantity 3 over x plus 2 over y ? the quantity 2y minus 3x all over the quantity 2y plus 3x the quantity 2y plus 3x all over the quantity 2y minus 3x the quantity 3y minus 2x all over the quantity 3y plus 2x the quantity 3y plus 2x all over the quantity 3y minus 2x I got C.

OpenStudy (shubhamsrg):

correct.

OpenStudy (anonymous):

mind helping me with another question?

OpenStudy (shubhamsrg):

hmm, what is it ?

OpenStudy (anonymous):

they all are correct

OpenStudy (anonymous):

What is the simplified form of the quantity x squared minus 5x plus 6 over 15 x y squared all over the quantity 2x squared minus 7x plus 3 over 5 x squared y ? its tricky to read idk how to make it not like that haha but these are the answer chices: 3y times the quantity 2x plus 1 over x times the quantity x plus 2 x times the quantity x plus 2 over 3y times the quantity 2x plus 1 x times the quantity x minus 2 over 3y times the quantity 2x minus 1 3y times the quantity 2x minus 1 over x times the quantity x minus 2

OpenStudy (anonymous):

that helps. lol @50_cent

OpenStudy (shubhamsrg):

your attempt being.. ?

OpenStudy (anonymous):

yes it was 40 questions and im stuck on ten questions and I have 7 now I go help from others as well!(:

OpenStudy (shubhamsrg):

I am asking whats your guess for this one ?

OpenStudy (anonymous):

ummmm im gonna go with A..

OpenStudy (shubhamsrg):

nop

OpenStudy (shubhamsrg):

re-try

OpenStudy (anonymous):

okay hold on..

OpenStudy (anonymous):

is it C? @shubhamsrg

OpenStudy (shubhamsrg):

bingo! ^^

OpenStudy (anonymous):

Thank you!

OpenStudy (shubhamsrg):

glad to help. Actually you don;t need much help, just a bit of confidence you need, you are doing the problems just fine. :)

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

\[\frac{ \frac{ 3 }{ x }-\frac{ 2 }{y } }{ \frac{ 3 }{x }+\frac{ 2 }{ y } }=\frac{ \frac{ 3y-2x }{ xy } }{\frac{ 3y+2x }{xy } } \]\[=\frac{ 3y-2x }{xy }\times \frac{ xy }{3y+2x }=\frac{ 3y-2x }{ 3y+2x }\]